A line has a slope of 7. Use the slope formula to find the missing coordinate
(x,8) (4,1)
x=5 but when I try to use the slope formula I am making a mistake somewhere. Please explain the steps you took to solve this problem.

Answer by stanbon(75887)

A line has a slope of 7. Use the slope formula to find the missing coordinate
(x,8) (4,1)
x=5 but when I try to use the slope formula I am making a mistake somewhere.
-----------------
If slope = 7 and (4,1) is a point,
the intercept is b where 1 = 7*4 + b; b = -27
---
Equation: y = 7x - 27
--------------------------
So, if y = 8, 8 = 7x - 27; 7x = 35; x = 5
============================================
Cheers,
Stan H.